Circular No. 6252

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1995ai IN NGC 7794 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery on

a Tech Pan film taken Oct. 23.94 UT of a supernova (V ∼ 18.3) at α = 23^{h}58^{m}33^{s}.06, δ = +10^{o}43′23′′.9 (equinox 2000.0), which is 15′′.7 west and 17′′.4 south of the center of NGC 7794 (and situated on a large, faint condensation of the southern arm of this spiral galaxy). M. Dennefeld, using the 1.2-m telescope at Observatoire de Haute Provence, estimated the flux (Cousins R band) of SN 1995ai on Oct. 24.78 UT to be ∼2 times fainter than the flux of each of two nearby stars of mag V ∼ 18.0 having position end figures 35^{s}.18, 51′′.1 and 36^{s}.42, 32′′.9. A third nearby star (V = 15.0) has position end figures 32^{s}.80, 42′59′′.7.

P. Garnavich, P. Challis, and R. Kirshner, Center for Astrophysics, write: “A CCD image obtained with the Mt. Hopkins 1.2-m telescope by N. Caldwell on Oct. 25.3 UT shows SN 1995ai at R = 17.9. A low signal-to-noise spectrogram taken by R. Jansen with the 1.5-m Tillinghast telescope shows a broad Hα feature with a P-Cyg profile, suggesting that this is a type-II supernova. Strong, narrow Hα, [N II], and [S II] lines at a redshifted velocity of 5100 km/s are superimposed on the spectrum.”

COMET C/1995 O1 (HALE-BOPP) A. Fitzsimmons, Queens University, Belfast, writes: “Optical spectra

(range 330–830 nm) of C/1995 O1 were obtained on Aug. 30.87 UT at the 4.2-m William Herschel Telescope (slit positioned at p.a. 357^{o}, with width 2′′.4 and length 44′′). Subsequent analysis has revealed the presence of the CN (0-0) 388-nm emission band with a flux of (1.7 ± 0.8) × 10^{−14} erg s^{−1} cm^{−2}. Assuming HCN as a parent, this corresponds to a production rate of Q(CN) ≈ 6 × 10^{25} molecules/s. No other emission features were observed.”

Total visual magnitude estimates: Sept. 28.06 UT, 10.3 (G. W. Kronk, Troy, IL, 0.33-m reflector); Oct. 1.11, 10.6 (R. Keen, Mt. Thorodin, CO, 0.32-m reflector); 13.00, 9.8 (D. W. E. Green, Harvard, MA, 0.25-m reflector); 22.12, 10.0 (C. S. Morris, Lockwood Valley, CA, 0.26-m reflector).

CORRIGENDA The Numbers of IAUCs 6244 and 6245, dated 1995 October 16 and 18,

respectively, were inadvertently cut off in their printing.

On IAUC 6247, line 11, for UGC 7122 read UGC 1722

1995 October 25 Daniel W. E. Green